5) A reaction has ΔH∘rxn= -118 kJ and ΔS∘rxn= 330 J/K .At what temperature is the change in entropy for the reaction equal to the change in entropy for the surroundings?

6) Given the values of ΔH∘rxn, ΔS∘rxn, and T below, determine ΔSuniv.
part a) ΔH∘rxn=− 90 kJ , ΔSrxn=− 150 J/K , T= 310 K
part b) ΔH∘rxn=− 90 kJ , ΔSrxn=− 150 J/K , T= 753 K
part c) ΔH∘rxn=+ 90 kJ , ΔSrxn=− 150 J/K , T= 310 K
part d) ΔH∘rxn=− 90 kJ , ΔSrxn=+ 150 J/K , T= 401 K
1
Expert's answer
2020-04-15T01:32:59-0400

5) We want to know the temperature at which:

-ΔH_rxn/T_surr = ΔS_rxn

Plugging in the numbers:

(118,000 J)/T_surr = (330 J/K)1p

T_surr = (118,000 J )/(330 J/K) = 357.6 K

6) a) ΔSsurr = -ΔH∘rx/T

ΔSsurr = +90 kJ/ 310 K = 0.2903 kJ/K = 290.3 J/K

ΔSuniv = ΔSsyst + ΔSsurr

ΔSuniv = - 150J/K + 290.3 J/K = 140.3 J/K

b) ΔSssur = +90 kJ/ 753 K = 0.1195 kJ/K = 119.5 J/K

ΔSuniv = - 150J/K + 119.5 J/K = 140.3 J/K

c) ΔSssur = - 90 kJ/ 310 K = - 290.3 J/K

ΔSuniv = - 150J/K + (-290.3) J/K = - 440.3 J/K

d) ΔSssur = +90 kJ/ 401 K = 0.2244 kJ/K = 224.4 J/K

ΔSuniv = 150J/K + 224.4 J/K = 374.4 J/K